OpenStudy (anonymous):

Find the probability of rolling a 4 or an even number on a number cube?

OpenStudy (aaliyahfairgood):

rolling a four is a 16.66666667% chance all even numbers is a 50% chance

OpenStudy (anonymous):

i need to write it as a fraction

OpenStudy (anonymous):

the numbers she just listed over 100

OpenStudy (anonymous):

okay im lost can someone explain

Miracrown (miracrown):

Well a number cube has six sides. So to roll a 4 or an even number we can roll: 2,4, or a 6 - So that gives us three possibilities.

OpenStudy (anonymous):

and thats 1/2 right ?

Miracrown (miracrown):

Yep, that's correct! :) We can roll: 2, 4 or 6. That is a total of 3 options. Total possibilities = 6. Probability = 3/6 = 1/2

OpenStudy (anonymous):

thank you (:

Miracrown (miracrown):

Yw (: